OpenStudy (anonymous):

True or False? If a*b=a*c, then either a=0 or b=c. The answer is false but I want to know how to get there with work. These are vectors, it has arrows above each letter and 0.

jimthompson5910 (jim_thompson5910):

assuming these are dot products, if a dot b = a dot c, then both dot products could be equal to zero ie a dot b = 0 a dot c = 0 which would mean that a dot b = a dot c is true

jimthompson5910 (jim_thompson5910):

this would mean that if 'a' is a vector that's perpendicular to both vectors b and c, then a dot b = a dot c is true but b = c doesn't necessarily have to be true and 'a' could be a nonzero vector

OpenStudy (anonymous):

a*b=a*c a*b-a*c=0 a*(b-c)=0 either a=0 orb-c=0, which gives a=0, or b=c a is parallel b-c or vector a is parallel to the plane containing the vectors b and c

OpenStudy (anonymous):

here*=cross product

OpenStudy (anonymous):

From a=0 or b=c, what's the next step to show that the statement is false?

jimthompson5910 (jim_thompson5910):

just pick any two vectors b, c where they aren't equal to each other then find the cross product of those vectors b and c to get the vector a You'll most likely find that 'a' is a nonzero vector

OpenStudy (anonymous):

So a doesn't =0 and b doesn't =c, right?

jimthompson5910 (jim_thompson5910):

yeah both are possible and you can still have a*b=a*c true (* is the dot product)

OpenStudy (anonymous):

Thanks.